OpenStudy (anonymous):

PR has an endpoint at (25, -5) and a midpoint of (18, -1). What is the value of the x-coordinate of the other endpoint?

OpenStudy (anonymous):

@anikhalder

OpenStudy (anikhalder):

Hmm...Okay..lets first draw a st. line...on the st line name the left pt P and the right pt Q and M the midpt.. @Muddingirl tell me when you have done that

OpenStudy (anonymous):

i did it @anikhalder

OpenStudy (anikhalder):

Now write the coord of pts:- Q (25,-5) M (18,-1) P (x,y) @Muddingirl ...tell me after you do this

OpenStudy (anonymous):

complete @anikhalder

OpenStudy (anikhalder):

Great... Now do you know the midpt formula...or the section formula?

OpenStudy (anonymous):

i know the midpoint formula @anikhalder

OpenStudy (anikhalder):

Awesome...:) pls tell me..:)

OpenStudy (anonymous):

m=(x1+x2/2,y1+y2/2)

OpenStudy (anikhalder):

Awesome..:) x1=x(P's x coord) x2=25(Q's x cord) M's x cord=18 and so M's x cord=x1+x2/2 Just place the values and solve for x1(or x)..:) @Muddingirl

OpenStudy (anonymous):

i get 11?

OpenStudy (anikhalder):

Excellent! You can similarly solve for y coord of P @Muddingirl

OpenStudy (anonymous):

oh really that's right?

OpenStudy (anikhalder):

11+25/2=36/2=18 ! that's the x coord of the midpt! :) You are right!

OpenStudy (anonymous):

oh wow, i actually solved right! I guess you were right practice makes perfect! Thanks

OpenStudy (anikhalder):

Welcome..keep practising!